the perimeter of a rectangle is 116m. the length is 7 m more than twice the width. find the dimensions
Answer by mananth(16946)

let width be x
length = 2x+7
...
Perimeter = 2L+2W
2x+2(2x+7)=116
..
2x+4x+14=116
6x+14=116
-14 -14
6x=116-14
6x=102
/6
x= 102/6
x= 17 m the width.
Length = 2*17+7 = 41 m
...
m.ananth@hotmail.ca